OpenStudy (rainbow_rocks03):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number greater than 1 is rolled? A. 167 B. 833 C. 667 D. 333

OpenStudy (rainbow_rocks03):

@kropot72 @NeonStrawsForever @TheSmartOne

OpenStudy (rainbow_rocks03):

No direct answers

OpenStudy (rainbow_rocks03):

TY

OpenStudy (anonymous):

Shame I'll take dat back

OpenStudy (rainbow_rocks03):

LOL it's alright. :)

OpenStudy (anonymous):

I tried to help

OpenStudy (rainbow_rocks03):

:) thx for trying

OpenStudy (kropot72):

The probability of rolling any number is 1/6. There are 5 numbers greater than 1. Therefore the probability of rolling a number greater than 1 is given by \[\large \frac{1}{6} \times\frac{5}{1}=\frac{5}{6}\] Now you need to multiply the probability of rolling a number greater than 1 by the number of rolls.

OpenStudy (rainbow_rocks03):

So I need to multiply 5/6 by something?

OpenStudy (kropot72):

Yes. Multiply the fraction by the given number of rolls.

OpenStudy (rainbow_rocks03):

oh so multiply 5/6*1000=833.333333333=833 so my answer is B.

OpenStudy (kropot72):

Correct.

OpenStudy (rainbow_rocks03):

YAY! thx for the help I will tag you when I need more help. :)

OpenStudy (kropot72):

You're welcome :)